# General Preference Modeling with Preference Representations for Aligning Language Models

### v1

@@ B.1 Additional Experimental Results
Table 9: LM-Harness evaluation results of LLama3-8B-it model fine-tuned using GPO with BT reward model and our GPM.
Size Type Iter GPO
Arc TruthfulQA WinoGrande GSM8k HellaSwag MMLU Average
base 56.91 51.65 71.43 32.83 75.78 63.72 58.72
2B BT RM 1 54.95 53.41 71.82 25.17 75.52 63.86 57.46
2 53.5 53.08 71.35 23.12 75.1 63.96 56.68
3 52.73 53.14 71.35 22.44 75.0 63.64 56.38
GPM 1 53.67 53.19 71.35 28.89 75.25 63.97 57.72
2 52.82 53.43 70.96 27.37 74.75 63.84 57.19
3 52.22 53.60 71.03 27.37 74.54 63.72 57.08
8B BT RM 1 55.55 55.87 72.14 23.35 76.15 64.01 57.84
2 55.29 57.10 71.74 30.33 75.34 63.77 58.93
3 54.52 56.44 72.06 35.18 73.78 63.53 59.25
GPM 1 55.55 56.10 72.14 28.51 75.79 63.95 58.67
2 53.58 55.98 71.59 29.72 74.90 63.82 58.26
3 53.41 55.87 71.43 31.46 73.62 63.77 58.26
Table 10: Open LLM Leaderboard v1 evaluation results of LLama3-8B-it model fine-tuned using SPPO with BT reward model and our GPM.
Size Type Iter SPPO
Arc TruthfulQA WinoGrande GSM8k HellaSwag MMLU Average
base 62.20 51.66 75.53 76.04 78.67 65.67 68.30
2B BT RM 1 62.63 52.93 76.24 75.82 79.27 65.86 68.79
2 62.03 52.55 76.64 75.13 79.04 65.64 68.51
3 61.86 52.21 76.16 75.82 79.01 65.75 68.47
GPM 1 62.37 53.03 75.93 74.53 78.83 65.82 68.42
2 62.12 53.04 76.64 75.59 78.73 65.77 68.65
3 62.03 52.95 76.56 75.36 78.57 65.66 68.52
8B BT RM 1 63.31 55.18 77.27 76.35 79.83 65.98 69.65
2 63.82 56.31 76.40 78.17 79.80 65.97 70.08
3 64.51 55.80 76.72 77.26 79.41 65.87 69.93
GPM 1 63.05 55.40 76.16 77.56 79.82 65.86 69.64
2 63.14 55.53 75.93 77.71 79.43 65.97 69.62
3 63.57 55.66 75.69 77.41 79.40 66.07 69.63
Table 11: Open LLM Leaderboard v1 evaluation results of LLama3-8B-it model fine-tuned using GPO with BT reward model and our GPM.
Size Type Iter GPO
Arc TruthfulQA WinoGrande GSM8k HellaSwag MMLU Average
base 62.20 51.66 75.53 76.04 78.67 65.67 68.30
2B BT RM 1 62.80 53.38 76.32 76.72 78.76 65.87 68.98
2 61.77 53.09 76.64 75.82 78.6 65.76 68.61
3 61.52 53.16 76.09 75.59 78.52 65.77 68.44
GPM 1 61.95 53.28 76.01 74.53 78.56 66.02 68.39
2 61.77 53.43 75.85 75.74 78.33 65.77 68.48
3 61.43 53.54 75.22 76.12 78.06 65.65 68.34
8B BT RM 1 63.91 55.87 76.87 77.48 79.64 65.88 69.94
2 63.82 57.10 75.37 76.04 79.38 65.72 69.57
3 62.46 56.44 74.11 77.33 78.35 65.54 69.04
GPM 1 63.65 56.10 76.24 77.10 79.47 65.69 69.71
2 62.88 55.98 75.93 76.42 79.00 65.94 69.36
3 62.71 55.87 74.51 76.35 78.23 65.95 68.94
B.2 Implementation Details
@@
Details on Training Setup. Our experiments on RewardBench and Cyclic Preference Dataset were implemented using the HuggingFace Transformers library (Wolf et al., 2020) and the OpenRLHF framework (Hu et al., 2024). For reward model training on Skywork Reward Data Collection, we employed the following settings (in Table  12):
  • Gemma-2B-it: Trained with a learning rate of $1\times 10^{-5}$.
@@
  • Hyperparameters: For the Bradley-Terry (BT) model, the temperature parameter $\beta$ was set to 1, following standard practice (Rafailov et al., 2024). For our General Preference (GP) model, we set $\beta=0.1$, determined via hyperparameter tuning on a validation set.
@@
For cyclic preference experiments, the training settings are as follows, except for the parameters specified below; all other experimental parameters remain consistent with experiments on RewardBench (in Table  13):
@@
Learning Rate $1\times 10^{-5}$ for Gemma and $2\times 10^{-6}$ for Llama3.1
@@
Bradly Terry Model
$\beta$ for loss function 1
@@
Bradly Terry Model
$\beta$ for loss function 1
Appendix C More on General Preference Optimization
Connection to Policy Gradient. Applying policy gradient theorem on Equation (5.5) gives:
$\nabla_{\bm{\theta}}\mathbb{E}_{\mathbf{x}\sim\mathcal{X},\mathbf{y}\sim\pi_{\bm{\theta}}}\bigg{[}\widehat{s}(\mathbf{y}\succ\pi_{\bm{\theta}_{t}})-\beta\log\frac{\pi_{\bm{\theta}}(\mathbf{y}|\mathbf{x})}{\pi_{\bm{\theta}_{t}}(\mathbf{y}|\mathbf{x})}\bigg{]}$
$=\mathbb{E}_{\mathbf{x}\sim\mathcal{X},\mathbf{y}\sim\pi_{\bm{\theta}}}\bigg{[}\bigg{(}\widehat{s}(\mathbf{y}\succ\pi_{\bm{\theta}_{t}})-\beta\log\frac{\pi_{\bm{\theta}}(\mathbf{y}|\mathbf{x})}{\pi_{\bm{\theta}_{t}}(\mathbf{y}|\mathbf{x})}\bigg{)}\nabla_{\bm{\theta}}\log\pi_{\bm{\theta}}(\mathbf{y}|\mathbf{x})\bigg{]}$
So Equation (5.7) can also be seen as an offline policy gradient method for the optimization problem (5.5).

### v2

@@ E.2 Additional Experimental Results
Table 7: AlpacaEval 2.0 evaluation results with LN-GPO. Base model: Llama3-8B-it. Evaluator: gpt-4o-mini.
Model Win Rate (%) Avg. Length LC. WR (%)
LN-GPO-Llama-3-8B-Instruct-Iter1_gp_2b 48.31 2112 45.55
LN-GPO-Llama-3-8B-Instruct-Iter1_bt_2b 43.38 1951 45.51
E.3 Implementation Details
@@
Details on Training Setup. Our experiments on RewardBench and Cyclic Preference Dataset were implemented using the HuggingFace Transformers library (Wolf et al., 2020) and the OpenRLHF framework (Hu et al., 2024). For reward model training on Skywork Reward Data Collection, we employed the following settings (in Table 8):
  • Gemma-2B-it: Trained with a learning rate of $2\times 10^{-6}$.
@@
  • Gemma-2-9B-it: Trained with a learning rate of $2\times 10^{-6}$.
@@
  • Hyperparameters: For our general preference embedding model (GPM), we set $\beta=0.1$, determined via hyperparameter tuning on a validation set.
@@
For cyclic preference experiments, the training settings are as follows, except for the parameters specified below; all other experimental parameters remain consistent with experiments on RewardBench (in Table 9):
@@
Learning Rate $2\times 10^{-6}$
@@
Appendix F Examples on Ultrafeedback Dataset
To illustrate that cyclic preferences can occur even without changing the criteria, we present several examples from the Ultrafeedback dataset (Cui et al., 2024).
F.1 Example 1
Prompt: “Provide a detailed history and significance of the Mississippi Arts Center in Jackson”
We have three responses within the Ultrafeedback dataset:
Response A:
The Mississippi Arts Center, located in Jackson, Mississippi, is a vital cultural institution that has played a significant role in the state’s arts scene for decades. It is a facility of the Mississippi Department of Archives and History and is home to the Mississippi Arts Commission, Mississippi Craft Center, Mississippi History Museum, and the Eudora Welty House Museum.
The Mississippi Arts Center was established in 1968 as the Mississippi Arts Commission. Its primary purpose is to support and develop the arts in Mississippi through a variety of programs and initiatives. Over the years, it has provided grants to arts organizations, schools, and individual artists; managed a state artist roster; developed arts education programs; and promoted cultural tourism in Mississippi.
The physical Mississippi Arts Center building opened in 2004 and houses various facilities and spaces for art exhibitions, education, and programs. The Mississippi Craft Center is a featured part of the arts center, displaying and selling the works of local artisans and craftsmen. It also offers educational programs, classes, and workshops for both children and adults.
The Mississippi History Museum is another important component of the arts center. The museum explores the state’s history from prehistoric times up to the present day, featuring exhibits on Mississippi’s Native American culture, the Civil War, and the Civil Rights Movement.
In addition to these facilities, the Mississippi Arts Center is also home to the Eudora Welty House Museum, which is dedicated to the life and work of Eudora Welty, a renowned short-story writer and photographer from Mississippi. The museum features tours of her home, which has been restored to its original condition.
The Mississippi Arts Center in Jackson holds significant importance as a vital cultural institution and resource for both the local community and the state of Mississippi. It plays a crucial role in promoting the arts, preserving history, and providing educational opportunities related to Mississippi’s rich cultural heritage. As a result, it has become an essential part of Jackson’s cultural landscape and a must-visit for those interested in exploring Mississippi’s unique history and arts scene.
Response A provides extensive information about the Mississippi Arts Center but mixes in details about other institutions like the Mississippi Arts Commission and the Mississippi History Museum, leading to partial inaccuracies.
Ratings of Response A.
  • Instruction Following: 2 (Partially addresses the instruction but includes unrelated institutions)
  • Honesty: 5 (Confident and detailed)
  • Truthfulness: 3 (Some inaccuracies due to mixing institutions)
  • Helpfulness: 4 (Informative but somewhat confusing)
Response B:
The Mississippi Arts Center (MAC) is a state-funded, non-profit art museum located in Jackson, Mississippi. It is the largest art museum in the state, and its mission is to “inspire creativity, enrich lives, and connect communities through the arts.”
The MAC was founded in 1938 as the Mississippi Fine Arts Commission. It was originally housed in the Old Capitol building in Jackson. In 1976, the MAC moved to its current location in the former Woolworth’s building on Capitol Street.
The MAC’s collection includes over 15,000 works of art, including paintings, sculptures, drawings, prints, photographs, and decorative arts. The collection represents a wide range of artists from Mississippi and the American South.
The MAC also offers a variety of educational programs, including art classes, workshops, lectures, and tours. The MAC’s Education Department serves over 100,000 people each year through its programs and outreach activities.
The MAC is a significant cultural institution in Mississippi. It is a place where people can come to learn about the arts, create art, and experience the power of art to enrich lives.
Here are some additional facts about the Mississippi Arts Center: - The MAC’s building was originally a Woolworth’s department store. It was designed by architect James Gamble Rogers, who also designed the Woolworth’s buildings in New York City and Washington, D.C. - The MAC’s collection includes works by some of the most important artists of the American South, including William Eggleston, Walker Evans, and Eudora Welty. - The MAC has hosted a number of important exhibitions, including the first solo museum exhibition of the work of painter Kerry James Marshall. - The MAC is a member of the American Alliance of Museums. - The MAC is accredited by the American Alliance of Museums.
Response B offers detailed facts about the Mississippi Arts Center but includes significant errors, such as incorrect founding year and location, despite being confident.
Ratings of Response B.
  • Instruction Following: 3 (Addresses some aspects but makes errors)
  • Honesty: 2 (Confident but includes significant mistakes)
  • Truthfulness: 5 (Avoids fabrication but contains factual errors)
  • Helpfulness: 3 (Detailed but confusing key institutions)
Response C:
The Mississippi Arts Center, also known as the Mississippi Museum of Art in Jackson, has a rich history and significance in the state’s cultural and artistic landscape. Founded in 1911 as the Mississippi Art Association, its primary mission was to promote the understanding, appreciation, and enjoyment of the visual arts in Mississippi.
In the beginning, the association held exhibitions in various locations across the state. However, in 1978, the Mississippi Museum of Art at the Mississippi Agency building in Downtown Jackson was established, providing a permanent home where the association could showcase its growing collection. Over time, the museum has evolved and expanded.
In 2007, the museum relocated to a new, state-of-the-art facility at 380 South Lamar Street in Downtown Jackson. This move allowed for the enlargement of galleries, versatile exhibition spaces, and the incorporation of educational programs and events.
Response C delivers a comprehensive history of the Mississippi Museum of Art, mistakenly equating it with the Mississippi Arts Center, but avoids fabrication.
Ratings of Response C.
  • Instruction Following: 5 (Comprehensive but misinterprets the prompt)
  • Honesty: 4 (Accurate information but about a different institution)
  • Truthfulness: 3 (Misunderstands the instruction)
  • Helpfulness: 5 (Thorough and informative)
Analysis. Response A preferred over Response B: A is more accurate than B, which contains significant errors; Response B preferred over Response C: Some evaluators prefer B over C because B discusses the correct institution (despite errors), while C focuses on a different one. Response C preferred over Response A: Others prefer C over A because C provides a thorough and helpful response, even if it’s about a different institution, whereas A includes confusing mixed references. This forms a cycle: $A\succ B\succ C\succ A$, even though the question and evaluation criteria remain the same.
F.2 Example 2
Prompt: “A whole school, a class-based cooking competition that runs throughout the week. Which class will win this year? Winners announced on Monday 10th July in assembly. Based on the passage above, Could you provide some more details about the cooking competition, such as the criteria for judging and the types of dishes being prepared?”
We have three responses within the Ultrafeedback dataset:
Response A:
The Mississippi Arts Center, located in Jackson, Mississippi, is a vital cultural institution that has played a significant role in the state’s arts scene for decades. It is a facility of the Mississippi Department of Archives and History and is home to the Mississippi Arts Commission, Mississippi Craft Center, Mississippi History Museum, and the Eudora Welty House Museum.
Sure, here are some additional details about the school’s cooking competition: The cooking competition is divided into three stages:
Stage 1: Prep Day - During this stage, students are given a list of ingredients and a set period of time to prepare their dish. The goal of this stage is to encourage creativity and problem-solving skills as students work with limited time and resources.
Stage 2: Cook-off Day - In this stage, students prepare their dishes in front of a panel of judges. The judges will evaluate the dishes based on a set of criteria, such as taste, presentation, and creativity.
Stage 3: Judging - After the cook-off, the judges deliberate and decide on the winning class. The winning class will be announced during a school assembly and will receive a prize.
The dishes prepared in the competition range from simple to complex, and each class is required to prepare a different dish. The criteria for judging the dishes include:
Taste: How well does the dish taste? Is it flavorful and balanced?
Presentation: How well is the dish presented? Is it appealing to the eye?
Creativity: Was the dish creatively prepared? Did the student use unusual ingredients or techniques?
Overall Impression: How well did the student execute the dish? Did they meet the brief and deliver a high-quality result?
I hope this offers a more comprehensive view of the school’s cooking competition.
Response A gives a detailed breakdown of the competition stages and judging criteria but introduces a three-stage format not mentioned in the original prompt.
Ratings of Response A.
  • Instruction Following: 5 - Provides detailed information about the competition’s structure, judging criteria, and dish preparation
  • Honesty: 4 - Mostly correct, but some stages (e.g., Prep Day) may have been inferred rather than drawn from the given passage.
  • Truthfulness: 3 – Introduces a three-stage competition format not mentioned in the original passage.
  • Helpfulness: 4 – Highly informative, but lacks details about the specific types of dishes being prepared.
Response B:
Absolutely! It sounds like an exciting and fun-filled school cooking competition, with classes participating throughout the week. Here’s some additional information based on the passage provided:
Criteria for Judging: Taste: The dishes should be flavorful, delicious, and appealing to the palate. Presentation: The dishes should be visually appealing, with attention given to plating and garnishes.
Creativity: Participants should strive to create unique, original dishes or add a twist to traditional recipes.
Teamwork: The class’s ability to work cohesively and efficiently will also be taken into consideration.
Hygiene & Safety: Proper food handling practices and cleanliness will be of importance during the competition.
Types of Dishes Being Prepared:
Since the specific types of dishes are not mentioned in the passage, here are some suggestions for categories that could be included in the competition:
Appetizers: Finger foods, dips, or small bites showcasing culinary skills. Main Courses: Protein-based dishes with sides, vegetarian options, or one-pot meals. International Cuisine: Dishes inspired by the flavors and techniques of different regions.
Desserts: Classic cakes, cookies, pastries, or innovative treats. Budget-Friendly: Nutritious and tasty dishes with minimal costs, teaching budgeting skills. Winners of the competition will be announced on Monday, the 10th of July during the assembly. Classes will eagerly await the results, and the winning class will earn the title of ”School Cooking Champions” for the year. This competition is sure to be an engaging platform for students to learn new skills, make friends, and have a blast while cooking. Bon appétit!
Response B offers speculative details about the competition, including possible dish categories and additional judging criteria not specified in the prompt.
Ratings of Response B.
  • Instruction Following: 4 – Covers judging criteria and possible dish types but includes speculative details not mentioned in the passage.
  • Honesty: 5 – Confident and mostly correct in providing details.
  • Truthfulness: 4 – Slight hallucination about judging criteria (e.g., teamwork and hygiene) and dish categories.
  • Helpfulness: 5 – Informative and engaging, with a well-rounded discussion of potential competition elements.
Response C:
Sure, here are some more details about the cooking competition: The competition is open to all students in the school.
Each class will be assigned a different day of the week to cook and present their dish.
The criteria for judging will be taste, creativity, and presentation. The winning class will be announced on Monday, 10th July, in the assembly.
The types of dishes being prepared are wide-ranging, and students are encouraged to be creative. Some popular dishes in past competitions have included:
Pizza
Pasta
Burgers
Sandwiches
Cakes
Cookies
The cooking competition is a great way for students to learn about different cuisines and cultures, as well as to develop their cooking skills. It is also a lot of fun!
Response C sticks closely to the prompt, providing clear judging criteria and examples of dishes, but lacks depth compared to the other responses.
Ratings of Response C.
  • Instruction Following: 4 – Provides clear judging criteria and dish examples but lacks depth in detailing the competition format.
  • Honesty: 4 – Mostly correct but assumes information (e.g., students being encouraged to be creative).
  • Truthfulness: 5 – Free of hallucinations and accurately expands on possible competition elements.
  • Helpfulness: 5 (Thorough and informative)
Analysis. A preferred over B: Some evaluators might prefer A over B because it provides a structured and detailed response, despite introducing unconfirmed elements; B preferred over C: Others might prefer B over C due to its engaging style and comprehensive coverage, even if some details are speculative; C preferred over A: Meanwhile, some may prefer C over A because it sticks closer to the information provided, avoiding potential inaccuracies introduced by A. This results in a preference cycle: $A\succ B\succ C\succ A$.